# Quantifying and Mitigating Privacy Risks for Tabular Generative Models

### v1

@@ 6. Evaluation Analysis
Table 3. Impact of DP-SGD training on DP-CTGAN, DP-TabDDPM, and the proposed DP-TLDM. Here, ”Resem.” stands for Resemblance, ”Distrim.” refers to Discriminability, ”S-out” denotes singling out attacks, and ”Link” represents linkability attacks.
Evaluation setup. Identical to Section 4, we use resemblance, discriminability and utility to measure synthetic data quality. We evaluate the privacy risk, ranging from 1-100, for four types of attacks, singling out, linkability, AIA and MIA. The privacy measure is the theoretical separation value in (4), representing the maximum difference of typeI-typeII error between the random guess and DP-SGD, illustrated in Fig 4. Three separation values are evaluated, namely $[0.1,0.15,0.2]$, where lower values indicate a stronger privacy level. We conduct the DP-SGP training on each generator under a given $\sigma$ value until the budget of separation depletes. For a fair comparison, we also apply batch clipping on all three synthesizers as outlined in (5).
@@
Figure 5. Comparison between three DP-enhanced synthesizers on various datasets.
6.1. Overview
We first present the overall performance across DP-CTGAN, DP-TabDDPM, and DP-TLDM in Table 3. The specific separation value is $0.1$, which is the most meaningful DP protection level in our evaluation. We summarize the key observations as follows.
@@ 6.1. Overview
DP protection yields a notable reduction in the risk of MIA on DP-TabDDPM and DP-TLDM. Notably, among all attacks considered, the most pronounced enhancement is observed in the context of MIA, where the risk diminishes substantially from approximately 90 to around 10. Given that MIA exploits additional information about the training dataset and model, its potential implications for the privacy of synthetic data are particularly severe. However, the DP mechanism employed here effectively mitigates these risks, successfully defending against MIA.
@@
6.2. Impact of privacy budget
@@
DP-TLDM consistently exhibits the best synthetic data quality across varying levels of privacy budget. Across all four datasets, a distinct hierarchy emerges among the three synthesizers, with DP-TLDM surpassing DP-TabDDPM and DP-CTGAN. This can be explained by the two benefits of our two-stage training scheme: firstly, Diffusion Models (DDPM) inherently exhibit great resilience to noisy input (Daras et al., 2024; Hsieh et al., 2023). By integrating the autoencoder with DP, the algorithm outputs latent representations with added perturbations. The resilience of the diffusion model ensures the generation of high-quality synthetic data. Secondly, the isolated two-stage training approach, where the privacy budget is solely allocated to the autoencoder stage, ensures that the diffusion process can refine and generate synthetic data without further compromising privacy. This efficient use of the privacy budget allows for the production of synthetic data that is not only of high quality but also adheres strictly to the required privacy constraints. The diffusion stage, not requiring additional privacy budget, acts as a compensatory mechanism for any potential decrease in data utility due to the privacy-preserving perturbations introduced in the autoencoder stage.
@@
Overall, these findings underscore that:
The two-stage training scheme of DP-TLDM, which leverages the inherent robustness of diffusion models to noisy inputs, achieves the optimal privacy-utility tradeoff among three DP-generators at equivalent privacy levels.
7. Conclusion
Motivated by the increasing adoption of synthetic tables as a privacy-preserving data sharing solution, we first conduct an empirical analysis on an extensive set of tabular generative models, addressing critical aspects of the privacy-utility trade-off through the lens of eight attacks, including singling out, linkability, AIA and five different MIAs. The experimental findings highlight distinctive performance characteristics across various generative models, with the tabular diffusion model demonstrating the highest data quality, albeit with notable privacy vulnerabilities against MIA. We then design DP-TLDM, a latent tabular diffusion trained by DP-SGD, following the $f$-DP framework. The key components of DP-TLDM are i) an autoencoder network to transform the tabular data into a compact and unified latent representation, and ii) a latent diffusion model to synthesize the latent tables. Thanks to the two-component design, and by applying DP-SGD to train the autoencoder, DP-TLDM obtains a rigorous DP guarantee, measured by the separation value. Our evaluation results against tabular GAN and regular tabular diffusion models trained with DP-SGD show that DP-TLDM can effectively mitigate the empirical privacy risks of synthetic data while achieving 15-50% higher data quality than other synthesizers with a stringent theoretical privacy budget.
Appendix A Nomenclature
@@
  • $\mathbb{D}$
    Domain of the randomized mechanism, representing the dataset space.
  • $\mathbb{R}$
    Range of the randomized mechanism, representing the output space.
  • $\mathcal{D}$
  • $\mathcal{E}$
@@
  • $\mathcal{N}(0,(C\sigma)^{2}\mathbf{I})$
  • $\mathcal{T}(\mathcal{M}(D),\mathcal{M}(D^{\prime}))(\alpha)$
@@
  • $\sigma$
  • $\tilde{g}_{r}$
  • $\tilde{X}$
    Reconstructed data from the latent representation by the decoder.
  • $\varepsilon$
  • $B$
    Batch size in the training process.
  • $b$
    Expected sample (mini-batch) size in DP-SGD.
  • $C$
    Gradient norm bound for clipping gradients during DP-SGD training.
  • $D$, $D^{\prime}$
    Neighboring datasets that differ by only a single record.
  • $E$
    Total number of training epochs in DP-SGD.
  • $E_{1},E_{2}$
  • $f$
  • $G_{\sigma^{-1}}$
@@
Appendix B Additional Results
B.1. Additional MIA Results
@@
Table 4. Memebership inference Attack additional results containing NaiveGroundhog (NG), HistGroundhog (HG), and Closest Distance-Hamming (CD-H), Closest Distance-L2 (CD-L) and Kernel Estimator (KE) attacks.
B.2. Additional DP Results
@@
Figure 6. Impact of varying $\sigma$ on DP for the Housing dataset with constant separation value, $sep=0.1$
Figure 7. Impact of varying $\sigma$ on DP for the Adult dataset with constant separation value, $sep=0.1$
Figure 8. Impact of varying $\sigma$ on DP for the Cardio dataset with constant separation value, $sep=0.1$

### v2

@@ 5 Performance Evaluation
Table 2: Impact of DP-SGD training on DP-CTGAN, DP-TabDDPM, and the proposed DP-TLDM. Here, “Resem.” stands for Resemblance, “Distrim.” refers to Discriminability, “S-out” denotes singling out attacks, and “Link” represents linkability attacks.
Evaluation setup. Identical to Section 0.C, we use resemblance, discriminability and utility to measure synthetic data quality. We evaluate the privacy risk, ranging from 1-100, for four types of attacks, singling out, linkability, AIA and MIA. The privacy measure is the theoretical separation value in (3), representing the maximum difference of typeI-typeII error between the random guess and DP-SGD, illustrated in Fig 3. Three separation values are evaluated, namely $[0.1,0.15,0.2]$, where lower values indicate a stronger privacy level. We conduct the DP-SGP training on each generator under a given $\sigma$ value until the budget of separation depletes. For a fair comparison, we also apply batch clipping on all three synthesizers as outlined in (4).
@@
Figure 4: Comparison between three DP-enhanced synthesizers on various datasets.
5.1 Overview
We first present the overall performance for DP-CTGAN, DP-TabDDPM, and DP-TLDM in Table 2. The specific separation value is $0.1$, which is the most meaningful DP protection level in our evaluation. We summarize the key observations as follows.
@@ 5.1 Overview
DP protection yields a notable reduction in the risk of MIA on DP-TabDDPM and DP-TLDM. Notably, among all attacks considered, the most pronounced enhancement is found in MIA, where the risk diminishes substantially from approximately 90 to around 10. Given that MIA exploits additional information about the training dataset and model, its potential implications for the privacy of synthetic data are particularly severe. However, DP mechanism employed here effectively mitigates these risks, successfully defending against MIA.
@@
5.2 Impact of privacy budget
@@
DP-TLDM consistently exhibits the best synthetic data quality across varying levels of privacy budget. Across all four datasets, a distinct hierarchy emerges among the three synthesizers, with DP-TLDM surpassing DP-TabDDPM and DP-CTGAN. This can be explained by the two benefits of our two-stage training scheme: firstly, Diffusion Models (DDPM) inherently exhibit great resilience to noisy input [10, 25]. By integrating the autoencoder with DP, the algorithm outputs latent representations with added perturbations. The resilience of the diffusion model ensures the generation of high-quality synthetic data. Secondly, the isolated two-stage training approach, where the privacy budget is solely allocated to the autoencoder stage, ensures that the diffusion process can refine and generate synthetic data without further compromising privacy. This efficient use of the privacy budget allows for the production of synthetic data that is not only of high quality but also adheres strictly to the required privacy constraints. The diffusion stage, not requiring additional privacy budget, acts as a compensatory mechanism for any potential decrease in data utility due to the privacy-preserving perturbations introduced in the autoencoder stage.
@@
Overall, these findings underscore that: The two-stage training scheme of DP-TLDM, which leverages the inherent robustness of diffusion models to noisy inputs, achieves the optimal privacy-utility tradeoff among three DP-generators at equivalent privacy levels.
6 Conclusion
Motivated by the increasing adoption of synthetic tables as a privacy-preserving data sharing solution,we design DP-TLDM, a latent tabular diffusion trained by DP-SGD, following the $f$-DP framework. Key components of DP-TLDM are i) an autoencoder network to transform tabular data into a compact and unified latent representation, and ii) a latent diffusion model to synthesize latent tables. Thanks to the two-component design, and by applying DP-SGD to train the autoencoder, DP-TLDM obtains a rigorous DP guarantee, measured by the separation value. Our evaluation results against tabular GAN and regular tabular diffusion models trained with DP-SGD show that DP-TLDM can effectively mitigate the empirical privacy risks of synthetic data while achieving 15-50% higher data quality than other synthesizers with a stringent theoretical privacy budget.
Acknowledgments
This research is part of the Priv-GSyn project, 200021E_229204 of Swiss National Science Foundation, and the DEPMAT project, P20-22 / N21022, of the research programme Perspectief which is partly financed by the Dutch Research Council (NWO).
Appendix 0.A Nomenclature
@@
$\alpha^{t}$
Parameters defining variance and transformation across timesteps.
$b$
Expected sample (mini-batch) size in DP-SGD.
$B$
Batch size in the training process.
$\beta^{t}$
Parameters defining variance and transformation across timesteps.
$C$
Gradient norm bound for clipping gradients during DP-SGD training.
$D$, $D^{\prime}$
Neighboring datasets that differ by only a single record.
$\mathcal{D}$
@@
$\mathbb{D}$
Domain of the randomized mechanism, representing the dataset space.
$E$
Total number of training epochs in DP-SGD.
$\mathcal{E}$
$E_{1},E_{2}$
@@
$\varepsilon$
$f$
$G_{\sigma^{-1}}$
$\tilde{g}_{r}$
@@
$p_{\theta}(x_{t-1}|x_{t})$
Reverse process, estimating data recovery from noise.
$q(x_{t}|x_{t-1})$
Forward process at step $t$, modeling transition probabilities.
@@
$\mathbb{R}$
Range of the randomized mechanism, representing the output space.
@@
$\tilde{X}$
Reconstructed data from the latent representation by the decoder.
@@
Appendix 0.B Risk-Utility Quantification
Figure 5: The risk-utility quantification and enhancing framework: from training to generation. Starting with the original tabular data, synthetic data is generated by synthesizers with or without DP protection. This synthetic data is then evaluated from two critical perspectives: utility (resemblance, discriminability, and utility) and privacy risk (singling out, linkability, attribute inference attack (AIA), and membership inference attack (MIA)).
0.B.1 Generative models
GAN-based models
We consider three different Generative Adversarial Networks (GAN) models. First, CTGAN [48] employs GANs with a focus on conditional generation. It employs mode-specific normalization to handle non-Gaussian distributions in continuous columns, and a conditional generator to address class imbalance in categorical columns. Second, CopulaGAN [41] improves upon CTGAN by utilizing cumulative distribution function-based transformations with Gaussian Copulas. It also performs inference using a likelihood approach, enhancing CTGAN’s ability to learn real data trends. Third, ADS-GAN [49] is a conditional GAN framework that generates synthetic data while minimizing re-identification risk. It achieves a certain degree of anonymization by incorporating a record-level identifiability metric into the generator’s loss function.
Statistical models
Here we consider Gaussian Copula models. In the Gaussian Copula (GC) method [34], the training data is used to obtain a Gaussian joint probability distribution that captures both marginal distributions and interdependence structures.
Difussion models
Diffusion models have recently become the leading paradigm in generative models for computer vision and NLP. In our framework, we consider TabDDPM [30], which extends diffusion models to tabular datasets, outperforming existing GAN/VAE alternatives. It employs the Gaussian diffusion process, a key component of the original DDPM [22], to model numerical columns effectively. It also uses the multinomial diffusion process to model categorical and binary features and to introduce uniform noise across classes to corrupt data.
0.B.2 Utility metrics
Resemblance
Discriminability
Utility
0.B.3 Threat model
In particular, for linkability attacks, the needed auxiliary data comprises two disjoint sets of attributes $T[:,A]$ and $T[:,B]$, extracted from $T$. For attribute inference attacks, the attacker needs to know the values of a set of attributes $T[:,A]$. In shadow modeling-based membership inference attacks, the attacker, alongside the target set $T$, has access to another reference dataset $X_{R}$. As in [40], this dataset mirrors the distribution of the training dataset and may or may not have overlapping records with it.
0.B.4 Attacks
As mentioned before, to quantify the privacy risks of tabular data synthesizers, we employ four distinct attacks in our evaluation framework. Three of these attacks—singling out, linkability, and attribute inference attack (AIA)—are derived from the guidelines set by the European General Data Protection Regulation (GDPR), following [19]. Additionally, we incorporate the membership inference attack (MIA), a well-established but previously omitted facet in [19]. This enhances our privacy quantification framework by allowing for the measurement of an additional and widely acknowledged dimension of privacy in synthetic data.
The implementation of singling out, linkability and AIA follows the open-source library Anonymeter  [^4] https://github.com/statice/anonymeter from [19] while the membership inference attack (MIA) follows the TAPAS toolbox  [^5] https://github.com/alan-turing-institute/tapas from [24].
Singling out
The singling out attack [19] aims to create predicates from the synthetic dataset that could identify an individual present in the training dataset. For example, if an attacker can determine that a dataset has only one individual with attributes like age: 25, height: 168, weight: 62, and cholesterol: 1, that individual is considered ”singled out”.
Following the implementation in [19], two algorithms are applied, the univariate algorithm and the multivariate algorithm. Both algorithms are based on the intuition that unique values or combinations of unique values in the synthetic data may also be unique in the original data. In the univariate algorithm, unique values are sampled for each attribute to obtain a random selection of predicates. In the multivariate algorithm, this is done for full records to obtain multivariate predicates.
Linkability
The linkability attack [19] aims to associate two or more records. The linkability arises when an attacker has two disjoint sets of attributes $X_{A}$ (e.g., age and height) and $X_{B}$ (e.g., weight and cholesterol level) from the original dataset, so that it can use the synthetic dataset to determine that two records $x_{a}\in X_{A}$ and $x_{b}\in X_{B}$ belong to the same individual.
Attribute inference attack (AIA)
The AIA attack [19] involves deducing undisclosed attribute values from information in the synthetic dataset. If an attacker knows certain attributes of an individual, such as age: 25, height: 168, and weight: 62, they might use the synthetic dataset to infer the cholesterol level of the individual.
Membership inference attack (MIA)
The MIA attack [38] aims to determine if a specific data record is present in the training dataset. MIA attacks have gained substantial attention within the research community, leading to various proposed strategies for inferring the membership status of synthetic data points. In our framework, we employ three types of MIA strategies, based on shadow modeling, distance and probability. This diverse set accommodates a range of adversarial scenarios, recognizing different adversary capabilities and constraints.
In the shadow modeling approach, given a reference dataset $X_{R}$, which shares the same distribution as the training dataset, and a target record $x_{t}$, multiple training sets $X_{i}$ from $X_{R}$ are sampled and shadow models trained to generate synthetic datasets $X_{s}$ from $X_{i}$ and $X_{s}^{\prime}$ from $X_{i}^{\prime}=X_{i}\cup x_{t}$. A classifier is then trained on labeled synthetic datasets $X_{s}$ and $X_{s}^{\prime}$ to predict the presence of target records in the training data $X_{train}$. To reduce the effect of high-dimensionality and sampling uncertainty, instead of directly training on $X_{s}$ and $X_{s}^{\prime}$, the NaiveGroundhog strategy uses a basic feature set $F_{naive}$ for training, while HistGroundhog utilizes a histogram feature set $F_{hist}$ with marginal frequency counts of each data attribute [40].
Metrics
Relative Risk Indicator The attacker success rate is a common metric for Membership Inference Attacks (MIA). However, for singling out, likability, and attribute inference attacks, as noted by [19], there’s a distinction: certain information might be inferred from patterns inherent in the entire population $X_{ori}$ rather than solely from the training dataset $X_{train}$ and its synthetic counterpart.
Following [19], the original dataset $X_{ori}$ is split into two disjoint partitions: $X_{train}$ and $X_{control}$. Privacy risk is then quantified by comparing attacker success rates for targets drawn from $X_{train}$ and $X_{control}$:
$R=\frac{\hat{\tau}_{train}-\hat{\tau}_{control}}{1-\hat{\tau}_{control}}$ (5)
Appendix 0.C Empirical Analysis
In this section, we put our risk-utility quantification framework described in Appendix B.3 to the test on publicly available datasets that have been extensively employed in tabular data analysis and synthesis.
0.C.1 Datasets
We employ four datasets, two small (up to 20000 samples) and two larger. Since small datasets usually make models prone to overfitting, by comparing these datasets, we can understand how dataset size and overfitting affect the quality and privacy of synthetic data. Some characteristics of the datasets are listed in Table 4 in the appendix.
The Loan dataset [3] contains demographic information on 5000 customers. It holds 14 features divided into 4 different measurement categories, including binary, interval, ordinal, and nominal features. The Housing dataset [43] relates to houses in a given California district and provides summary statistics based on the 1990 Census data. It comprises 20,640 instances with 1 categorical and 9 numerical features and a total of 207 missing values. The Adult dataset [4] contains information on individuals’ annual incomes and related variables. It consists of 48842 instances with 14 mixed datatype features in total, and a total of 6465 missing values. The Cardiovascular Heart Disease dataset [11] contains detailed information on the risk factors for cardiovascular disease, including 70000 instances with 13 mixed-type columns.
0.C.2 Privacy-utility Trade-off
Table 3 presents detailed results quantifying both utility and risk aspects of synthetic data for all four datasets employing the five generative models described in Appendix B.1. We present the three utility metrics discussed before (i.e., resemblance, discriminability, and utility), where a higher score indicates better performance, as well as the privacy risk for the four attacks considered (Singling out, Linkability, AIA, MIA), where a lower risk indicates better performance. Due to space reasons, we keep the detailed statistics of the five MIA attacks in the appendix.
Dataset Method Quality Score $\uparrow$ Privacy Risk $\downarrow$
Resem. Discri. Utility S-out Link AIA MIA
Loan CopulaGAN 92 95 70 52.81 2.31 6.98 2.86
CTGAN 92 85 93 54.90 0.00 8.11 2.86
ADS-GAN 93 95 73 17.23 0.00 0.00 22.86
GC 86 82 78 28.68 0.00 0.00 5.72
TabDDPM 98 100 97 26.31 2.23 16.68 45.72
Housing CopulaGAN 94 90 62 8.14 0.00 1.54 20.00
CTGAN 94 92 64 12.55 0.45 0.00 20.00
ADS-GAN 93 87 74 1.73 1.43 0.98 48.58
GC 91 84 32 4.16 0.00 0.00 5.72
TabDDPM 96 98 93 1.30 0.16 0.00 88.58
Adult CopulaGAN 93 97 81 10.25 0.05 5.08 17.14
CTGAN 90 79 83 20.18 0.55 3.16 10.00
ADS-GAN 88 59 83 19.74 0.00 0.00 20.00
GC 80 50 56 32.80 0.38 2.86 8.58
TabDDPM 96 98 98 22.72 0.46 0.00 94.28
Cardio CopulaGAN 87 93 96 66.54 1.13 28.75 0.00
CTGAN 84 68 97 62.04 1.11 24.07 11.42
ADS-GAN 90 71 100 59.76 0.89 15.21 31.42
GC 81 63 86 61.02 0.44 6.30 22.86
TabDDPM 95 99 100 60.77 1.31 23.08 94.28
Table 3: Quantification of Risk-Utility for Five Generative Models Across Various Datasets. Here, “Resem.” stands for Resemblance,“Distrim.” refers to Discriminability, “S-out” denotes singling out attacks, and ”Link” represents linkability attacks.
Regarding MIA strategies, notable effectiveness is achieved by the NaiveGroundhog (NG), HistGroundhog (HG), and Closest Distance-Hamming (CD-H) strategies, which are able to reach success rates of 60% or higher in some cases. These results are detailed in Table 5 in Appendix D.1. Remarkably, HistGroundhog consistently outperforms other MIA strategies when applied to the TabDDPM synthesizer. In contrast, the NaiveGroundhog and Closest Distance-Hamming strategies demonstrate better efficacy when employed on other synthesizers.
This leads us to conclude that in our experiments:
No. Rows No. Columns Missing Values
Loan 5000 14 0
Housing 20640 10 0.1%
Adult 48842 14 0.94%
Cardio 70000 13 0
Table 4: Characteristics of four tabular datasets used in our study.
Appendix 0.D Additional Results
0.D.1 Additional MIA Results
@@
Table 5: Memebership inference Attack additional results containing NaiveGroundhog (NG), HistGroundhog (HG), and Closest Distance-Hamming (CD-H), Closest Distance-L2 (CD-L) and Kernel Estimator (KE) attacks.
0.D.2 Additional DP Results
@@
Figure 6: Impact of varying $\sigma$ on DP for the Housing dataset with constant separation value, $sep=0.1$
Figure 7: Impact of varying $\sigma$ on DP for the Adult dataset with constant separation value, $sep=0.1$
Figure 8: Impact of varying $\sigma$ on DP for the Cardio dataset with constant separation value, $sep=0.1$